OpenStudy (jozelynw):

Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model.

OpenStudy (jozelynw):

How to make a extraneous solution?

OpenStudy (jozelynw):

@ikram002p

OpenStudy (jozelynw):

@welshfella

OpenStudy (jozelynw):

do you know how to write a extraneous radical equation

OpenStudy (xapproachesinfinity):

extraneous sol is solution that is not in the domain of the equation you started with

OpenStudy (jozelynw):

huh

OpenStudy (jozelynw):

well, then how do I create a radical extraneous equation?

OpenStudy (xapproachesinfinity):

consider the following equation \(\sqrt{\frac{x+1}{x}}=\sqrt{\frac{1}{x}}\) see that x=0 is not in the domain try to solve that equation

OpenStudy (xapproachesinfinity):

it is easier to create an equation with correct solution (not extraneous )

OpenStudy (welshfella):

-3 sqrt(7 + 2x) = 24 sqrt(7 + 2x) = -8 square both sides 7 + 2x = 64 2x = 57 x = 57/2 now plug in x = 57/2 and see if it satisfies the original equation

OpenStudy (xapproachesinfinity):

consider the following equation \(\sqrt{x+1}=\sqrt{3x-1}\) solution is \(x=1\) if we check it on the equation we it checks out

OpenStudy (jozelynw):

it wants me to use this equation a√x+b+c=d

OpenStudy (xapproachesinfinity):

you just said radical equation you did specify any details

OpenStudy (xapproachesinfinity):

well at any rate check @welshfella work

OpenStudy (jozelynw):

I know because then you would just be giving me answers and then people would be thinking that im cheating

OpenStudy (welshfella):

a negative number times a radical will give an extraneous solution

OpenStudy (jozelynw):

so i just wanted to know how to create a radical extraneous equation

OpenStudy (jozelynw):

can you give me an example

OpenStudy (welshfella):

if you plug in 57/2 into the above it won't give you 24

OpenStudy (jozelynw):

but it's suppose to

OpenStudy (jozelynw):

28.5

OpenStudy (welshfella):

no - ones like that will give an extraneous solution When you square you get a positive result so a negative value can get 'lost'.

OpenStudy (jozelynw):

so 28.5 is not 57/2

OpenStudy (welshfella):

-3 * sqrt(7 + 28.5) = -17.87

OpenStudy (jozelynw):

what do you mean sqrt ? sqrt of 3

OpenStudy (welshfella):

do you see that this is an extraneous solution?

OpenStudy (welshfella):

square root of (7 + 28.5)

OpenStudy (jozelynw):

i do it equals -17.87

OpenStudy (jozelynw):

wait did something wrong

OpenStudy (jozelynw):

sqrt of 35.5 =5.95 and 5.95 times -3 equals -17.87

OpenStudy (welshfella):

yes that is correct so 28.5 is not a solution its extraneous

OpenStudy (welshfella):

now one that does not have an extraneous solution would be sqrt(4x + 1) = 7 so 4x + 1 = 49 4x = 48 x = 12 which will fit the original equation

OpenStudy (welshfella):

here is a link which deals with the subject http://www.purplemath.com/modules/solverad2.htm

OpenStudy (jozelynw):

what are the options

OpenStudy (welshfella):

what do you mean?

OpenStudy (jozelynw):

you said " which will fit the original equation"

OpenStudy (welshfella):

I meant if you replace x by 12 in the original equation then the equation will balance the left side will equal 7

OpenStudy (welshfella):

sqrt(4*12 + 1) = sqrt 49 = 7

OpenStudy (welshfella):

so x = 12 is a valid solution

OpenStudy (jozelynw):

so the rule to this is.......

OpenStudy (welshfella):

a negative value as above will give an extraneous solution. But there' more to it than that that's why i posted the link. Some equations will have one valid and one extraneous solution , for example.

OpenStudy (welshfella):

sorry but i have to go now

OpenStudy (jozelynw):

thank you so much for you help @welshfella

OpenStudy (welshfella):

yw